OpenStudy (anonymous):

find the volume of the square pyramid. length = 12 cm, width = 12 cm height =11 cm. I think it's 792..

OpenStudy (anonymous):

@Michele_Laino please help!

OpenStudy (michele_laino):

the requested volume is: \[V = \frac{{L \times W \times H}}{3} = \frac{{12 \times 12 \times 11}}{3} = ...?\]

OpenStudy (anonymous):

528?

OpenStudy (michele_laino):

that's right! 

OpenStudy (michele_laino):